Suppose that the terms of trade between Mexico and the US are 3.5 tons of avocados in exchange for 1 ton of soybeans. Would the two countries be willing to trade at this relative price? Explain.

Production Alternatives (Tons) US Production Possibilities

Product

R

S

T

U

V

Avocados

0

30

33

60

90

Soybeans

30

20

19

10

0

Production Alternatives (Tons) Mexico’s Production Possibilities

Product

A

B

C

D

E

Avocados

0

20

24

40

60

Soybeans

15

10

9

5

0
Answer #1

Opportunity cost of 1 ton of Soyabean in US = Total avocados given up/Total Soyabeans produced = 90/30 = 3 avocados

Opportunity cost of 1 ton of Soyabean in Mexico = Total avocados given up/Total Soyabeans produced = 60/15 = 4 avocados

They would be willing to trade as long as trading ratio lies between 3 and 4 avocados. So, the two countries be willing to trade at the relative price of 3.5 tons of avocados in exchange for 1 ton of soybeans.